8/x=2/5
x=
can you explain how to do it and also the answer I save this info to study
x/2=5/15
x=

Answer by Alan3354(69443)

8/x=2/5
-------------
The shortcut is to "cross multiply."
2x = 8*5
2x = 40
x = 20
---------